OpenStudy (lilai3):

WILL AWARD MEDAL/FAN! How can you find the number of sides of an equiangular polygon by measuring one of its interior angles? By measuring one of its exterior angles? Thank you for your help! It is greatly appreciated!~

OpenStudy (anonymous):

I really don't like dealing with the majority of geometry students here, but you're lucky. This is a straightforward question. Are you familiar with equiangular polygons?

OpenStudy (lilai3):

No, not really. I'm not really sure what they are. Can you please explain or draw it for me?

OpenStudy (anonymous):

According to traditional euclidean theory, equiangular polygons are figures that are at complete equilibrium. For example, a square would be an equiangular polygon as all its angles and sides are equal.

OpenStudy (anonymous):

The interior of any shape is derived from a circle. According to Pythagoras, a circle can technically (for this purpose) be defined as having an angle of "360". If you had a square, all of it's angles would be 90. 90*4=360

OpenStudy (anonymous):

Same for a triangle. If you cut a square in half, it's equivalent to cutting a circle in half. If you were able to tear and shape that triangle back into a circle, you would only have an angle of 180.

OpenStudy (anonymous):

Do you understand?

OpenStudy (lilai3):

oh, so does that mean a rectangle can also be an equiangular polygon? also a rhombus?

OpenStudy (lilai3):

@happyyanee4 ?

OpenStudy (e.mccormick):

A rhombus is only equiangular of it is a square.

OpenStudy (e.mccormick):

Do you know the calculation of the angles as determined by the number of sides?

OpenStudy (e.mccormick):

happyyanee4 gave you two clues to find the calculation. If sides are 4, then 4 * 90 = 360 If sides are 3, then 3 * 60 = 180 So both the angle and the number of sides change at the same time.

OpenStudy (lilai3):

so the answer is you have to know the number of sides first?

OpenStudy (e.mccormick):

No.

OpenStudy (lilai3):

i'm confused.

OpenStudy (e.mccormick):

Well, you know a 3 sided opbject happens to have a total of 180 degrees and that an equiangular version is 60 per angle. You likewise know that a 4 sided object happns to contain 360 segrees, but each angle is 90. So no matter ehat the total is, the individual angles in an equiangular object will be some number and that number changes.

OpenStudy (lilai3):

Wait so how do i describe the answer?

OpenStudy (e.mccormick):

Well, you have to describe that. And you have to know how to do it in relationship to both the internal and external angles.

OpenStudy (lilai3):

so you just say like "you have to know the degree of an angle"...?

OpenStudy (lilai3):

i don't know how to explain that relationship between the angle measure and the number of sides

OpenStudy (e.mccormick):

Well, there is a formula, which I asked if you knew. This one: http://www.regentsprep.org/regents/math/geometry/gg3/lpoly2.htm

OpenStudy (lilai3):

because if it was the other way around, you could find the measure with the number of sides, instead of finding the number of sides with a measure.

OpenStudy (lilai3):

yes, that was the one i was talking about ^ yes, i am familiar with it

OpenStudy (lilai3):

wait so basically, the answer to this problem is just the reverse of that formula?

OpenStudy (e.mccormick):

basically you need to revise it to find the number of sides.

OpenStudy (lilai3):

the reverse?

OpenStudy (e.mccormick):

Well, there is going to be a similar relationship. How much work you need to put into the answer depends on if they want a formula or an idea. Oh, and here is a neat one... you don't have to do much for the second question. 

OpenStudy (lilai3):

the second question's answer is: by knowing the interior angle?

OpenStudy (e.mccormick):

basically, yah. If you answer the first one, you can just describe the relationship of the exterior to interior angles. Then reference the first answer.

OpenStudy (lilai3):

yeah. the only problem? LOL we don't know the answer to the first one. well, that is, not really.

OpenStudy (e.mccormick):

Well, 3s=60, and 4s=90.... lets see if those solve to the same number.

OpenStudy (lilai3):

do you mean proportions?

OpenStudy (e.mccormick):

20 and ... nope... so something is missing from the easy method.

OpenStudy (lilai3):

no, they don't.

OpenStudy (lilai3):

yeah

OpenStudy (e.mccormick):

Hmm... but that formula used n-2... so that will be involved somehow. And I bet if you looked you could find the formula for finding the n of a n gon from the angle. Just trying to cook one up will take a little more work.

OpenStudy (lilai3):

180 (n-2) = n ^ formula for finding n? ^

OpenStudy (e.mccormick):

Yah, probably easiest to just solve that for n.

OpenStudy (lilai3):

but n could be anything, since n is an unknown polygon, right?

OpenStudy (e.mccormick):

But you would need the sum of interiaor angles... hmmm.... so might still have another thing to solve.

OpenStudy (lilai3):

i'm sorry but i'm so lost right now. like the first question's answer is just use the formula to figure out the number of sides, right?

OpenStudy (e.mccormick):

No. Because you don't know the sum.

OpenStudy (lilai3):

oh yeah....oops

OpenStudy (e.mccormick):

Hmmm... geometric sequence.

OpenStudy (lilai3):

then how do we solve this problem?

OpenStudy (lilai3):

what do we need to do with the formula?

OpenStudy (e.mccormick):

Well, using the formula I found a 5 sided... 540 degrees... but that just tells me it is a geometric sequence. Not sure where happy was going with the circle.

OpenStudy (lilai3):

what?

OpenStudy (e.mccormick):

What is I am thinking rather than googling.

OpenStudy (lilai3):

wait me? i'm not googling anything LOL does it sound like i'm not trying? it's probably cuz i'm hecka tired. it's like three in the morn here. sorry. but i'm trying, so that i don't waste your time and mine. but math is just sooo not my subject.

OpenStudy (e.mccormick):

No, no. I am just thinking out loud rather than googling. So what I type is incomplete.

OpenStudy (lilai3):

oh psh.

OpenStudy (e.mccormick):

See, while the angles are getting larger, the difference btween each angle and the previous one is getting smaller, so not a simple sequence. A fraction in there...

OpenStudy (e.mccormick):

Hmmm... what if I did not work with interior? Exterior?

OpenStudy (e.mccormick):

OK, stuff I was playing with on a scrach pad... s=180(n-2) s/180=(n-2) s/180 +2=n 3 180 60 + 30 4 360 90 + 18 5 540 108 + 12 6 720 120 + 7 900 not a nice number... a_n - a_1 * r^{n-1} 180/n^2 270/n^2 Sides Ex-Ang 3 120 4 90 5 72 6 120 Most of it is junk, but I just noticed something. The 3 sided and 6 sided. 3 sides has an interior of 60 and exterior of 120. 6 sides has an interior of 120 and exterior of 60. To me, that suggests there is a numeric relationship.

OpenStudy (lilai3):

but other sides, like 5 and 7 don't really have a relationship...

OpenStudy (e.mccormick):

yah, but I know I am looking for a fraction, and it would be 5 and 10 to look at.

OpenStudy (e.mccormick):

So if I try interiors... x/3 = 60 x/6 = 120 Ewww... that won't work. Exteriors: x/3 = 120 x/6 = 60 3*120 = 360 6*60 = 360 Oooh.... And there is that circle happy talked about!

OpenStudy (e.mccormick):

So try a few more with the exterior angle. Divide 360 by the exterior angle and see if you get the number of sides.

OpenStudy (lilai3):

oh i see sort of

OpenStudy (e.mccormick):

8 sides = 1080 degrees, or 135 per angle, which is 45 on the exterior. That matches up nicely with the 90 on a 4 sided. =) 360/90 = 4 360/45 = 8

OpenStudy (e.mccormick):

So the key is what happy said, a circle as 360 degrees, and what the question asked about, interior and \(\textbf{exterior}\) angles.

OpenStudy (lilai3):

okay, thank you so much!!~

OpenStudy (lilai3):

you helped greatly! it's four now, and i must be getting some sleep. thank you!